### 12 Mar 2026 v1

# Why urban heterogeneity limits the 15-minute city

Abstract
The “15-minute city” has emerged as a central paradigm in urban planning, promoting universal access to work and essential services within short travel times. Its feasibility—particularly for commuting to work—has however rarely been examined quantitatively. Here, we show that proximity to employment is fundamentally constrained by the internal structure of urban economies. Combining urban geometry with empirically observed firm-size distributions, we derive a lower bound on commuting times that holds independently of planning choices or transport technologies. This bound reveals a sharp transition: when employment is sufficiently concentrated, no spatial rearrangement of workplaces can ensure uniformly short commutes, even under optimal placement. Applied to Paris and its near suburbs, we find that achieving universal 15-minute commutes would require substantial economic restructuring or differentiated mobility strategies. The relevant question is therefore not whether an $x$-minute city is achievable, but what the minimal feasible $x$ is given a city’s economic structure and spatial scale.
Main
Building on earlier traditions of neighborhood-based urbanism and accessibility planning, the “15-minute city” Moreno et al. (2021) (or more generally the $x$-minute city) promotes the idea that essential urban functions—work, commerce, healthcare, and education—should be reachable within a short travel time from one’s residence, ideally by walking or cycling. Initially framed as a response to climate change, social inequalities, and car-dependent urban growth, it has rapidly gained visibility as a normative model for sustainable urban development Moreno et al. (2025), promising reduced car dependence and emissions alongside improvements in quality of life, public health, and social inclusion Saelens et al. (2003).
Although employment is explicitly included in the original formulation Moreno et al. (2021), it has proven difficult to operationalize empirically, and commuting-to-work constraints have received limited quantitative attention. Concerns about job–housing proximity are longstanding: classic work on accessibility and job–housing balance already showed that residences in job-rich areas are associated with shorter commutes Levinson (1998). The 15-minute city can thus be viewed as a strengthening of this principle, aiming not merely at average reductions in commuting time but at universal proximity for all workers. Recent empirical work, however, shows that proximity to services does not necessarily translate into localized travel, particularly for work trips Zhang et al. (2025).
Methodological reviews confirm this gap between conceptual ambition and empirical practice, showing that the 15-minute city lacks a unified operational definition, relies on heterogeneous indicators, and predominantly measures access to local services, while commuting responds more weakly to proximity-based interventions Büttner et al. (2024); Omwamba et al. (2025); Papadopoulos et al. (2023); Pozoukidou and Chatziyiannaki (2021); Khavarian-Garmsir et al. (2023); Abbiasov et al. (2024); Rhoads et al. (2023). Large-scale empirical assessments further quantify substantial heterogeneity of accessibility within and across cities worldwide, highlighting the role of population density and resource distribution in shaping proximity outcomes Büttner et al. (2024); Omwamba et al. (2025); Bruno et al. (2024). Despite these advances, the feasibility of the $x$-minute city remains poorly understood at a more fundamental level. In particular, recent reviews Büttner et al. (2024); Omwamba et al. (2025) make clear that most studies stop short of addressing a structural question: whether universal proximity to work is achievable in principle, independently of specific planning choices, service redistribution strategies, or mobility technologies.
In particular, Bruno et al. Bruno et al. (2024) propose a global framework to quantify accessibility to essential services and to simulate their optimal redistribution across cities worldwide. They measure proximity times to nine categories of point-of-interests (outdoor activities, learning, supplies, eating, moving, cultural activities, physical exercise, services, and healthcare), and develop an algorithm to relocate these amenities so as to equalise services per capita and increase the fraction of the population living within a 15-minute condition. Their results highlight strong heterogeneity across cities and show that substantial service redistribution would be required in many urban contexts. However, the most significant structural obstacle to achieving a universal 15-minute city lies in commuting to work. Unlike local amenities such as shops or schools, employment is intrinsically heterogeneous: a small number of large firms employ a substantial fraction of the workforce, while many small establishments coexist alongside them. Crucially, employment is not a freely replicable service that can be redistributed without altering the underlying economic structure itself. This difference is therefore not merely quantitative but structural.
More precisely, firm sizes follow broad, approximately Zipfian distributions across countries, sectors, and time periods Ramsden and Kiss-Haypál (2000); Axtell (2001); Gaffeo et al. (2003); Gabaix (2016); Santos and Cajueiro (2024). Although firm-level employment data are typically available only in aggregated form at the urban scale, establishment counts reveal substantial within-city heterogeneity, with large employers accounting for a disproportionate share of employment (see section I of the SI). In general, we can expect that large metropolitan areas necessarily host the upper tail of the firm-size distribution Sridhar and Wan (2010), making economic heterogeneity a structural constraint at the urban scale.
This observation raises a fundamental question left open by existing studies Büttner et al. (2024); Omwamba et al. (2025): under what conditions is it possible, even in principle, to locate firms in space such that every worker can reach their workplace within a prescribed travel time $\tau_{0}$? We address this question using a minimal spatial framework combining three ingredients: (i) a population distributed over a city of area $A$, (ii) firms whose sizes follow a Zipf distribution with exponent $\gamma$, and (iii) an optimization problem in which firm locations are chosen to minimize the maximum commuting distance. This framework yields strict theoretical bounds on proximity, which we confront with numerical optimization and a case study of Paris and its near suburbs.
Model
To isolate the role of the journey-to-work commute, we introduce a minimal model that focuses on this single ingredient and reveals the existence of a structural lower bound on commuting time. We thus consider an idealized city of characteristic radius $R$ (typically of order $\sqrt{A/\pi}$, where $A$ is the urban area), and total employment $E$. Residents are assumed to be uniformly distributed in space unless stated otherwise. Employment is provided by $N$ firms whose sizes are heterogeneous and ranked by decreasing employment $m_{1}>m_{2}>\dots>m_{N}$. Firm sizes are assumed to follow a Zipf distribution, such that the firm of rank $r$ employs $m_{r}\propto r^{-\gamma}$ workers, where the exponent $\gamma$ controls the degree of economic inequality: $\gamma=0$ corresponds to homogeneous firms, while larger values of $\gamma$ indicate increasingly concentrated employment in a small number of large firms.
The framework is intentionally minimal and aims to identify feasibility constraints that are independent of detailed urban design, land-use regulation, or transport technologies. It applies most directly at spatial scales where employment locations compete for access to the same pool of residents, such as neighborhoods, districts, or central urban areas. At larger metropolitan scales, many cities exhibit a polycentric organization with multiple employment centers rather than a single dominant core Anas et al. (1998); Berroir et al. (2011); Bertaud and Malpezzi (2003); Le Néchet (2012); Louail et al. (2014). In such cases, proximity constraints should be evaluated at the level of individual subcenters rather than across the entire metropolitan area. Applying a 15-minute constraint at the metropolitan scale is therefore not meaningful in general, as commuting distances necessarily exceed this threshold even under optimistic mobility assumptions. Thus, proximity-based planning implicitly assumes a decomposition of cities into smaller functional units, and the present framework provides a feasibility test for such units.
The normalization condition $\sum_{r=1}^{N}m_{r}=E$, together with $m_{r}=m_{1}r^{-\gamma}$, fixes the size of the largest firm as
$m_{1}=\frac{E}{H_{N}(\gamma)},$ (1)
where $H_{N}(\gamma)=\sum_{r=1}^{N}r^{-\gamma}$ is the $n$-th generalized harmonic number. For large $N$, $H_{N}(\gamma)\sim N^{1-\gamma}/(1-\gamma)$ for $\gamma<1$, $H_{N}(1)\sim\log N$ in the marginal case, and $H_{N}(\gamma)\to\zeta(\gamma)$ for $\gamma>1$, where $\zeta(\gamma)$ denotes the Riemann zeta function. This result Eq. 1 shows that the firm-size distribution alone determines how many workers must be served by the largest employer, independently of spatial considerations.
Each worker is assigned to a single firm, and commuting is approximated by straight-line (Euclidean) distances, a standard first-order approximation in urban spatial models Batty (2013), which captures the underlying spatial constraint independently of network details. For a given spatial arrangement of firms, we define $\ell_{\max}$ as the maximum commuting distance across the population. The central feasibility question is whether firms can be located in space such that $\ell_{\max}\leq L_{0}$, where $L_{0}$ denotes a prescribed maximum commuting distance (or, equivalently, a maximum commuting time). Geometrically, the worker–firm assignment problem is equivalent to a weighted Voronoi partition Okabe et al. (2009), in which firm sizes act as weights constraining the spatial extent of their catchment areas. As firm-size heterogeneity increases, these weighted cells become increasingly distorted, forcing some regions to span large distances and driving the proximity feasibility transition.
Results
Theoretical feasibility constraints
We derive a strict theoretical constraint on the feasibility of an $x$-minute city that holds independently of any numerical optimization. This constraint follows from elementary geometric considerations combined with firm-size heterogeneity and defines an absolute lower bound that no spatial rearrangement of workplaces can overcome.
The main idea is that the dominant constraint is imposed by the largest firm. Assuming that firm sizes follow a Zipf distribution with exponent $\gamma$, the largest firm employs a fraction of the total workforce given by Eq. 1. Even if this firm is optimally located at the point of maximum population density, it must recruit its workforce from a finite catchment area which has a certain size. For a uniform worker density $\rho=E/A$, the radius $\ell_{\max}$ of this catchment area satisfies $\rho\pi\ell_{\max}^{2}=m_{1}$. Requiring that all workers commute less than a prescribed distance $L_{0}$ therefore yields the lower bound
$L_{0}\;\geq\;L_{0}^{\ast}\equiv\sqrt{\frac{A/\pi}{H_{N}(\gamma)}}.$ (2)
In the numerical model on a disk of radius $R$, this reduces to $L_{0}/R\geq 1/\sqrt{H_{N}(\gamma)}$. Expressed in terms of time rather than distance, the minimal commuting time achievable in this city is then
$\tau_{0}(\gamma)\;=\;\frac{1}{v}\,\frac{\sqrt{A/\pi}}{\sqrt{H_{N}(\gamma)}},$ (3)
where $v$ is the characteristic travel speed (walking or biking). Equation (3) constitutes a fundamental limit on urban proximity: it cannot be improved by redistributing firms, modifying transport infrastructure, or altering land-use patterns. It provides the reference against which numerical optimizations and empirical case studies are interpreted below (more details and discussion can be found in section II of the SI).
The factor $\sqrt{A/\pi}/v$ corresponds to the time required to traverse the city radius, while the effect of firm-size heterogeneity enters through the denominator $\sqrt{H_{N}(\gamma)}$. The resulting constraint depends qualitatively on the value of $\gamma$. For $\gamma>1$, the harmonic number converges to $H_{N}(\gamma)\to\zeta(\gamma)$ in the large-$N$ limit, yielding a finite asymptotic lower bound $L_{0}^{\ast}=\sqrt{A/\pi}/\sqrt{\zeta(\gamma)}$. In this regime, increasing the number of firms cannot reduce commuting distances below a fixed fraction of the city size, revealing a structural obstruction to universal short commutes. By contrast, in the homogeneous limit $\gamma=0$, one has $H_{N}(0)=N$ and the bound reduces to the purely geometric scaling $L_{0}^{\ast}=\sqrt{A/\pi}/\sqrt{N}$, which can be made arbitrarily small by increasing the number of firms.
The transition between these regimes reflects a qualitative change in the nature of proximity constraints, from geometry-dominated at low heterogeneity to extreme-value-dominated at large $\gamma$. Empirically, estimates based on binned establishment-size data for U.S. metropolitan areas yield values in the range $\gamma\in[0.4,1.0]$, with mean $\langle\gamma\rangle\simeq 0.71\pm 0.12$ (see the section I of the SI). For the Paris metropolitan area, available data are consistent with a larger exponent, $\gamma\simeq 1.4$ (see section V of the SI), placing it deep in the high-heterogeneity regime where proximity constraints become intrinsically difficult to satisfy at the metropolitan scale.
Numerical phase diagram
We next examine how the theoretical feasibility bounds manifest at finite system size using explicit spatial optimization. The aim of this numerical analysis is not to calibrate real cities, but to probe the robustness and sharpness of the feasibility constraint with finite numbers of workers and firms.
Figure 1: Comparison before and after optimization. Illustration of the simulated annealing optimization for a representative example ($E=1000$, $N=60$, $\gamma=1.5$, $m_{1}=424.6$, $m_{N}=1$). Firm locations are initially assigned at random (symbol size proportional to firm size), and workers are randomly allocated to firms. The initial configuration yields a maximal commuting distance $\ell_{\max}\approx 1.69$. After simulated annealing ($T_{\mathrm{init}}=0.5$, $T_{\min}=10^{-4}$, $\alpha=0.95$, $n_{\mathrm{steps}}=1000$ per temperature), the system converges to an optimized spatial arrangement with $\ell_{\max}\approx 0.70$, corresponding to a reduction of $\sim 60\%$. In the optimized configuration, the largest firms are located closer to the center of the system, which reduces extreme commuting distances and prevents the occurrence of very long trips.
Figure 2: Finite-size phase diagram of proximity feasibility. Optimized maximum commuting distance $\ell_{\max}$ as a function of the firm-size exponent $\gamma$, obtained from simulated annealing for a system with total employment $E=5000$, city radius $R=1$, establishment density $N=\alpha E$ with $\alpha=0.06$, and capacity tolerance $\delta=20\%$. Symbols show averages over $10$ independent spatial configurations. The dashed line indicates the theoretical lower bound $L_{0}^{\ast}$ given by Eq. 2 for a uniform population density. Green and red shaded regions denote, respectively, feasible and infeasible proximity regimes: below the bound, no spatial arrangement of establishments can achieve a smaller maximum commuting distance, whereas above it realizable configurations exist.
For fixed values of $(\gamma,E,N)$ and a prescribed capacity tolerance $\delta$, firm locations are optimized using simulated annealing Kirkpatrick et al. (1983) to minimize the maximum commuting distance $\ell_{\max}$. The tolerance $\delta$ quantifies the allowed mismatch between firm capacity and the number of workers assigned to it: a firm of size $m$ can accommodate a number of workers in the interval $(1-\delta)m,(1+\delta)m]$. This parameter captures realistic flexibility in firm capacity, congestion effects, and imperfect matching between workers and establishments. As derived in Section III of the SI, capacity flexibility modifies the theoretical lower bound by replacing the factor $1/H_{N}(\gamma)$ with $(1-\delta)/H_{N}(\gamma)$, yielding a reduced minimum commuting time $\tau_{\min}(\gamma,\delta)=\tau_{\min}(\gamma,0)\,\sqrt{1-\delta}$. The optimization is repeated over multiple independent realizations of the population distribution to estimate both the mean optimal $\ell_{\max}$ and its fluctuations (for more details on the numerical simulation, see the section IV in the SI). Throughout this analysis, we assume that the number of establishments scales proportionally with total employment, $N\simeq\alpha E$, following empirical observations for U.S. Metropolitan Statistical Areas, with $\alpha\simeq 0.06$. This assumption fixes the average firm size, $\langle m\rangle=1/\alpha$, and closes the model (see the SI).
We show in Fig. 1 the result of the optimization for a representative example with $E=1000$ employees and $N=60$ firms, with firm sizes distributed according to a Zipf law with exponent $\gamma=1.5$. Starting from a random spatial configuration of firms and a random allocation of workers to firms, the simulated annealing algorithm searches for the arrangement that minimizes the maximal commuting distance $\ell_{\max}$. In the initial configuration, the maximal commute is $\ell_{\max}=1.68$, which is close to the diameter $2$ of the disk. After optimization, the maximal commuting distance decreases to $\ell_{\max}\approx 0.70$, corresponding to roughly one quarter of the system diameter. The optimized configuration exhibits a clear spatial organization: the largest firm is located near the center of the system, while smaller firms are distributed toward the periphery. This arrangement minimizes extreme commuting distances and prevents the occurrence of very long trips.
The previous example illustrates how spatial optimization reduces extreme commuting distances in a single configuration. We now examine how this limit depends on the degree of economic heterogeneity. Figure 2 shows the optimized maximal commuting distance $\ell_{\max}$ as a function of the firm-size exponent $\gamma$, together with the theoretical lower bound $L_{0}^{\ast}$. Two main results emerge. First, the numerical solutions reproduce the qualitative trend predicted by the theory: increasing firm-size heterogeneity leads to a monotonic increase in the minimal achievable commuting distance. The transition from feasible to infeasible proximity regimes is sharp, confirming that heterogeneity acts as the dominant control parameter. Second, the optimized configurations lie systematically above the theoretical bound. This offset reflects finite-size effects and the discrete nature of the spatial optimization problem. While the analytical results provide asymptotic lower bounds valid in the limit of large populations, numerical realizations involve finite numbers of firms and residents. These finite-size constraints prevent the numerical solutions from reaching the theoretical bound, creating a systematic gap above it. The theoretical curve therefore acts as a strict lower bound, while the numerical results represent realizable configurations at moderate system size.
The discrepancy between numerical and theoretical bounds is largest around and above $\gamma\simeq 1$, a range consistent with empirically observed Zipf-like firm-size distributions. This value marks a qualitative crossover between weakly and strongly heterogeneous regimes. For $\gamma>1$, firm sizes span several orders of magnitude without a characteristic scale, and the optimization becomes dominated by a small number of extreme firms. In this regime, $\ell_{\max}$ is controlled by rare, system-spanning constraints rather than by typical firms, leading to an abrupt breakdown of proximity.
Overall, these numerical results show that the infeasibility transition predicted by the theory is already visible at moderate system sizes. In finite systems, proximity constraints are even stronger: the optimized maximum commuting distance is systematically larger than the theoretical lower bound.
Paris as a stress test for the 15-minute city
Paris provides a particularly challenging case for the feasibility of a 15-minute city. Approximately $1.8$ million individuals work inside the city of Paris, but commuting flows are highly asymmetric: nearly $60\%$ (1,080,0001{,}080{,}000 individuals) of workers employed in Paris reside outside the city (data for 2017; INSEE (2022)). Among these $1.08$ million workers, about $55\%$ come from the near suburbs (the petite couronne), with the remainder residing in more distant suburbs and other regions. We therefore focus on the $1.314$ million individuals who work in Paris and live either in Paris or in its near suburbs, excluding the outer suburbs (the grande couronne), where commuting distances typically exceed $20$ km and a 15-minute constraint is unrealistic.
Paris hosts approximately 406 000406\,000 establishments Greffe du Tribunal de Commerce de Paris (2018), corresponding to an average firm size of $\langle m\rangle\simeq 4.4$ employees. Rather than relying on a single estimate of $N$ for the combined area of Paris and the petite couronne (which for example would be $\sim 3\times 10^{5}$ for $4.4$ employees per firm), we vary $N$ over almost two orders of magnitude, from $10^{4}$ to $5\times 10^{5}$, encompassing any plausible value. To characterize firm-size heterogeneity, we estimate the Zipf exponent $\gamma$ using establishment-size data available at the scale of the Paris urban area. Using INSEE data for 2023 INSEE (2023), we obtain a relatively large value $\gamma\simeq 1.38$, indicating strong employment concentration. Given uncertainties in the estimation, we treat $\gamma$ as a control parameter and explore a broad range of values.
The central question is whether these $N$ (with $N\in[10^{4},5\times 10^{5}]$) establishments can be spatially reorganized over Paris and the petite couronne such that all $1.314$ million workers commute less than 15 minutes. To address this question, we use the theoretical lower bound $\tau_{\min}(\gamma)$ derived above. At this scale, population density can be reasonably approximated as uniform, allowing direct application of the analytical framework.
Figure 3 shows the resulting minimal commuting time $\tau_{0}(\gamma)$ as a function of firm-size heterogeneity for both walking and cycling, across a wide range of establishment counts $N$. Two main regimes emerge. For $\gamma\gtrsim 1$, the curves at different $N$ collapse onto a single line: in this regime $H_{N}(\gamma)$ is well approximated by $\zeta(\gamma)$, and the theoretical bound depends only on the area and the exponent, not on the number of firms. For $\gamma\lesssim 1$, by contrast, $H_{N}(\gamma)$ grows without bound with $N$, making $\tau_{0}$ sensitive to the precise establishment count; larger $N$ spreads workers across more, smaller catchment areas and thereby reduces the worst-case commuting distance. for Paris, the empirical firm-size distribution, however, falls in the $\gamma>1$ regime, where the bound is $N$-independent. The shaded bands quantify the additional gain from capacity flexibility $\delta$. The multiplicative reduction $\sqrt{1-\delta}$ is independent of $\gamma$ (see SI), but even for $\delta=0.3$—corresponding to a roughly $16\%$ decrease in $\tau_{0}$—the correction is far too modest to bring commuting times below the 15-minute threshold for any empirically relevant value of $\gamma$. Crucially, these results show that for a wide range of $\gamma$ values compatible with empirical estimates, including $\hat{\gamma}\simeq 1.38$, the 15-minute threshold cannot be reached—neither by increasing the number of establishments, nor by allowing substantial capacity flexibility. While increasing travel speed through cycling significantly reduces commuting times, it is generally insufficient to offset the constraint imposed by firm-size heterogeneity. These results indicate that, at the scale of Paris and its near suburbs, achieving a 15-minute city would require not only improvements in mobility but a substantial reorganisation of economic activity itself.
Figure 3: Theoretical minimum commuting time for Paris and its near suburbs. Minimal achievable commuting time $\tau_{0}(\gamma)$ as a function of the firm-size exponent $\gamma$, computed for the combined area of Paris and the petite couronne ($A=762~\mathrm{km}^{2}$). Solid curves correspond to walking ($v=5$ km/h) and dashed curves to cycling ($v=15$ km/h); colour encodes the number of establishments $N_{\rm est}$, with values $10^{3}$, $10^{4}$, $10^{5}$, $3\times 10^{5}$, and $5\times 10^{5}$. The shaded band beneath each curve shows the reduction in $\tau_{0}$ afforded by a capacity flexibility $\delta$ ranging from 0 (upper edge) to $0.3$ (lower edge), corresponding to a multiplicative factor $\sqrt{1-\delta}$. The horizontal dotted line marks the 15-minute threshold, while the vertical dashed line indicates the empirical estimate $\hat{\gamma}\simeq 1.38$.
Structural limits to urban proximity
Our results reveal a fundamental limitation of proximity-based urban ideals that arises from economic organization rather than from urban form or transport infrastructure. The key mechanism limiting proximity discussed here is firm-size heterogeneity. When employment is broadly distributed and includes very large establishments, a small number of firms must draw workers from spatially extended catchment areas to satisfy their capacity. These dominant employers therefore impose long commuting distances, even under optimal spatial placement. By contrast, when firms are relatively homogeneous and remain small, employment can be organized through many local catchment areas, allowing commuting distances to remain short throughout the city.
Firm-size heterogeneity is a robust empirical feature of modern economies and is likely to remain significant within cities, although more detailed urban-scale data are needed to quantify it precisely. If the Paris metropolitan area is representative, many large cities may lie deep in the heterogeneous regime identified here, where the theoretical lower bound on commuting time is already large. In this regime, short commutes for all workers cannot, in principle, be ensured solely through spatial reorganization, even under optimistic assumptions about travel speeds. This implies that the 15-minute city is not universally achievable, but depends on the specific economic and spatial characteristics of each city. While firm-size heterogeneity is a central ingredient of the present analysis, it is not the only factor shaping commuting patterns and accessibility. The purpose of this work is not to provide an exhaustive description of urban complexity, but to identify fundamental theoretical constraints that apply independently of particular planning choices.
Importantly, the framework developed here also offers a practical way to estimate the minimum achievable proximity time in a given city. When empirical data are available for the firm-size distribution (through $\gamma$, $N$, and $E$) and for the spatial extent $A$ of the urban area, the theory yields a lower bound on commuting time that no spatial rearrangement of workplaces alone can overcome. In this sense, the relevant question is not whether a city can be made into an “$x$-minute city” for a prescribed value of $x$, but rather what the smallest feasible $x$ is, given its economic structure and geometry. This perspective suggests a differentiated strategy: proximity-based planning may successfully ensure local access to daily amenities, while access to employment in heterogeneous urban economies may require complementary mobility solutions and efficient transport networks rather than purely spatial redistribution. By quantifying these structural limits, this work provides a framework for evaluating where and when proximity-based urban planning can realistically succeed.
Acknowledgments: I thank Jieun Lee and Chris Webster for interesting discussions on the 15 minute city.
Data availability statement: The datasets used in this study are freely available from public repositories: U.S. Census Bureau (2023); INSEE (2022); Greffe du Tribunal de Commerce de Paris (2018); INSEE (2023).
Competing interests: None.

### 16 Mar 2026 v2

@@
 (2025): under what conditions is it possible, even in principle, to locate firms in space such that every worker can reach their workplace within a prescribed travel time $\tau_{0}$?
-We address this question using a minimal spatial framework combining three ingredients: (i) a population distributed over a city of area $A$, (ii) firms whose sizes follow a Zipf distribution with exponent $\gamma$, and (iii) an optimization problem in which firm locations are chosen to minimize the maximum commuting distance.
+We address this question using a minimal spatial framework combining three ingredients: (i) a population distributed over a city of area AA, (ii) firms whose sizes follow a Zipf distribution with exponent $\gamma$, and (iii) an optimization problem in which firm locations are chosen to minimize the maximum commuting distance.
 This framework yields strict theoretical bounds on proximity, which we confront with numerical optimization and a case study of Paris and its near suburbs.
@@
 To isolate the role of the journey-to-work commute, we introduce a minimal model that focuses on this single ingredient and reveals the existence of a structural lower bound on commuting time.
-We thus consider an idealized city of characteristic radius $R$ (typically of order $\sqrt{A/\pi}$, where $A$ is the urban area), and total employment $E$.
+We thus consider an idealized city of characteristic radius $R$ (typically of order $\sqrt{A/\pi}$, where AA is the urban area), and total employment $E$.
 Residents are assumed to be uniformly distributed in space unless stated otherwise.
-Employment is provided by $N$ firms whose sizes are heterogeneous and ranked by decreasing employment $m_{1}>m_{2}>\dots>m_{N}$.
+Employment is provided by NN firms whose sizes are heterogeneous and ranked by decreasing employment $m_{1}>m_{2}>\dots>m_{N}$.
 Firm sizes are assumed to follow a Zipf distribution, such that the firm of rank $r$ employs $m_{r}\propto r^{-\gamma}$ workers, where the exponent $\gamma$ controls the degree of economic inequality: $\gamma=0$ corresponds to homogeneous firms, while larger values of $\gamma$ indicate increasingly concentrated employment in a small number of large firms.
@@
 where $H_{N}(\gamma)=\sum_{r=1}^{N}r^{-\gamma}$ is the $n$-th generalized harmonic number.
-For large $N$, $H_{N}(\gamma)\sim N^{1-\gamma}/(1-\gamma)$ for $\gamma<1$, $H_{N}(1)\sim\log N$ in the marginal case, and $H_{N}(\gamma)\to\zeta(\gamma)$ for $\gamma>1$, where $\zeta(\gamma)$ denotes the Riemann zeta function.
+For large NN, $H_{N}(\gamma)\sim N^{1-\gamma}/(1-\gamma)$ for $\gamma<1$, $H_{N}(1)\sim\log N$ in the marginal case, and $H_{N}(\gamma)\to\zeta(\gamma)$ for $\gamma>1$, where $\zeta(\gamma)$ denotes the Riemann zeta function.
 This result Eq.
@@
 The resulting constraint depends qualitatively on the value of $\gamma$.
-For $\gamma>1$, the harmonic number converges to $H_{N}(\gamma)\to\zeta(\gamma)$ in the large-$N$ limit, yielding a finite asymptotic lower bound $L_{0}^{\ast}=\sqrt{A/\pi}/\sqrt{\zeta(\gamma)}$.
+For $\gamma>1$, the harmonic number converges to $H_{N}(\gamma)\to\zeta(\gamma)$ in the large-NN limit, yielding a finite asymptotic lower bound $L_{0}^{\ast}=\sqrt{A/\pi}/\sqrt{\zeta(\gamma)}$.
 In this regime, increasing the number of firms cannot reduce commuting distances below a fixed fraction of the city size, revealing a structural obstruction to universal short commutes.
@@
 Starting from a random spatial configuration of firms and a random allocation of workers to firms, the simulated annealing algorithm searches for the arrangement that minimizes the maximal commuting distance $\ell_{\max}$.
-In the initial configuration, the maximal commute is $\ell_{\max}=1.68$, which is close to the diameter $2$ of the disk.
+In the initial configuration, the maximal commute is $\ell_{\max}=1.68$, which is close to the diameter 22 of the disk.
 After optimization, the maximal commuting distance decreases to $\ell_{\max}\approx 0.70$, corresponding to roughly one quarter of the system diameter.
@@
 Among these $1.08$ million workers, about $55\%$ come from the near suburbs (the petite couronne), with the remainder residing in more distant suburbs and other regions.
-We therefore focus on the $1.314$ million individuals who work in Paris and live either in Paris or in its near suburbs, excluding the outer suburbs (the grande couronne), where commuting distances typically exceed $20$ km and a 15-minute constraint is unrealistic.
+We therefore focus on the $1.314$ million individuals who work in Paris and live either in Paris or in its near suburbs, excluding the outer suburbs (the grande couronne), where commuting distances typically exceed 2020 km and a 15-minute constraint is unrealistic.
 Paris hosts approximately 406 000406\,000 establishments Greffe du Tribunal de Commerce de Paris (2018), corresponding to an average firm size of $\langle m\rangle\simeq 4.4$ employees.
-Rather than relying on a single estimate of $N$ for the combined area of Paris and the petite couronne (which for example would be $\sim 3\times 10^{5}$ for $4.4$ employees per firm), we vary $N$ over almost two orders of magnitude, from $10^{4}$ to $5\times 10^{5}$, encompassing any plausible value.
+Rather than relying on a single estimate of NN for the combined area of Paris and the petite couronne (which for example would be $\sim 3\times 10^{5}$ for $4.4$ employees per firm), we vary NN over almost two orders of magnitude, from $10^{4}$ to $5\times 10^{5}$, encompassing any plausible value.
 To characterize firm-size heterogeneity, we estimate the Zipf exponent $\gamma$ using establishment-size data available at the scale of the Paris urban area.
@@
 Given uncertainties in the estimation, we treat $\gamma$ as a control parameter and explore a broad range of values.
-The central question is whether these $N$ (with $N\in[10^{4},5\times 10^{5}]$) establishments can be spatially reorganized over Paris and the petite couronne such that all $1.314$ million workers commute less than 15 minutes.
+The central question is whether these NN (with $N\in[10^{4},5\times 10^{5}]$) establishments can be spatially reorganized over Paris and the petite couronne such that all $1.314$ million workers commute less than 15 minutes.
 To address this question, we use the theoretical lower bound $\tau_{\min}(\gamma)$ derived above.
 At this scale, population density can be reasonably approximated as uniform, allowing direct application of the analytical framework.
-Figure 3 shows the resulting minimal commuting time $\tau_{0}(\gamma)$ as a function of firm-size heterogeneity for both walking and cycling, across a wide range of establishment counts $N$.
+Figure 3 shows the resulting minimal commuting time $\tau_{0}(\gamma)$ as a function of firm-size heterogeneity for both walking and cycling, across a wide range of establishment counts NN.
 Two main regimes emerge.
-For $\gamma\gtrsim 1$, the curves at different $N$ collapse onto a single line: in this regime $H_{N}(\gamma)$ is well approximated by $\zeta(\gamma)$, and the theoretical bound depends only on the area and the exponent, not on the number of firms.
-For $\gamma\lesssim 1$, by contrast, $H_{N}(\gamma)$ grows without bound with $N$, making $\tau_{0}$ sensitive to the precise establishment count; larger $N$ spreads workers across more, smaller catchment areas and thereby reduces the worst-case commuting distance.
-for Paris, the empirical firm-size distribution, however, falls in the $\gamma>1$ regime, where the bound is $N$-independent.
+For $\gamma\gtrsim 1$, the curves at different NN collapse onto a single line: in this regime $H_{N}(\gamma)$ is well approximated by $\zeta(\gamma)$, and the theoretical bound depends only on the area and the exponent, not on the number of firms.
+For $\gamma\lesssim 1$, by contrast, $H_{N}(\gamma)$ grows without bound with NN, making $\tau_{0}$ sensitive to the precise establishment count; larger NN spreads workers across more, smaller catchment areas and thereby reduces the worst-case commuting distance.
+for Paris, the empirical firm-size distribution, however, falls in the $\gamma>1$ regime, where the bound is NN-independent.
 The shaded bands quantify the additional gain from capacity flexibility $\delta$.
@@
 Importantly, the framework developed here also offers a practical way to estimate the minimum achievable proximity time in a given city.
-When empirical data are available for the firm-size distribution (through $\gamma$, $N$, and $E$) and for the spatial extent $A$ of the urban area, the theory yields a lower bound on commuting time that no spatial rearrangement of workplaces alone can overcome.
+When empirical data are available for the firm-size distribution (through $\gamma$, NN, and $E$) and for the spatial extent AA of the urban area, the theory yields a lower bound on commuting time that no spatial rearrangement of workplaces alone can overcome.
 In this sense, the relevant question is not whether a city can be made into an “$x$-minute city” for a prescribed value of $x$, but rather what the smallest feasible $x$ is, given its economic structure and geometry.